OpenStudy (anonymous):

solve for h q=2h-2u What is the solution for h Write this phrase as a mathematical expression 6 more than 4 times x

OpenStudy (anonymous):

\[q=2h-2u\] add 2u to both sides \[q+2u=2h\] divide by 2 \[\frac{q+2u}{2}=h\]

OpenStudy (anonymous):

six more than four times x \[4x+6\]

OpenStudy (anonymous):

1. (q + 2u)/2 2. 4x + 6